Circular No. 5676

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J0422+32 T. Kato, S. Mineshige, and R. Hirata, Kyoto University, write: “CCD

photometry of the optical counterpart of GRO J0422+32 was obtained with the 0.6-m telescope at Ouda Station, Kyoto University, on five nights during Nov. 21-25. The analysis of the data reveals periodic modulation with a period of 0.217 ± 0.001 day and an amplitude of 0.10 mag in V. The light curve exhibits a steeper rise than decline and resembles superhumps in SU UMa-type stars. The secondary humps are not detected. The considerable increase in amplitude and the distinct light curve profile from that reported by Chevalier and Ilovaisky (IAUC 5644) suggest that the present modulation may have developed after their observations. If this modulation is caused by the tidal instability of the accretion disk, as in SU UMa stars, and with an assumption of the Roche lobe-filling main sequence secondary, the lower limit of the mass of the compact object is 2.2 M_{⊙}, which exceeds the 1.4 M_{⊙} of the canonical neutron star (Mineshige et al. 1992, Publ. Astron. Soc. Japan, 44, L15-18).”

GRO J1837+59 The Compton Observatory EGRET Team (Y. C. Lin, J. Chiang, J.

Fierro, P. F. Michelson, and P. L. Nolan, Stanford University; K. Brazier, G. Kanbach, H. A. Mayer-Hasselwander, and C. von Montigny, Max-Planck-Institut für Extraterrestrische Physik; D. L. Bertsch, B. L. Dingus, C. E. Fichtel, R. C. Hartman, S. D. Hunter, P. W. Kwok, J. R. Mattox, P. Sreekumar, and D. J. Thompson, Goddard Space Flight Center, NASA; D. A. Kniffen, Hampden-Sydney College; and E. Schneid, Grumman Aerospace Corporation) communicates: “Quick-look analysis of EGRET high-energy γ-ray data for the current Compton Observatory pointing shows a bright source (designated GRO J1837+59) coincident with one seen in the observation of 1991 May 30-June 8. The best estimate of the source position is l = 88^{o}.7, b = +24^{o}.8, with a 68-percent-confidence error radius of 0^{o}.4. The positional uncertainty is relatively large due to the large angle of the source from the detector axis. No catalogued sources at other wavelengths are obvious counterparts of the γ-ray source. The present Compton Observatory viewing continues until Dec. 22, and a more sensitive observation of this source region is scheduled for 1993 Mar. 9-23.”

1992 December 16 Daniel W. E. Green